solve the following differential equations

1. (1 + 2 𝑒 𝑥 𝑦) + 2𝑒 𝑥 𝑦 (1 − 𝑥 𝑦 ) 𝑑𝑦 𝑑𝑥 = 0 

2.1. 𝑡𝑎𝑛 𝑥 . 𝑠𝑖𝑛2 𝑦 𝑑𝑥 + 𝑐𝑜𝑠2 𝑥 . 𝑐𝑜𝑡 𝑦 𝑑𝑦 = 0

2. (1 + 2 𝑒 𝑥 𝑦) + 2𝑒 𝑥 𝑦 (1 − 𝑥 𝑦 ) 𝑑𝑦 𝑑𝑥 = 0

3. (𝑥 − 2𝑦 + 1) 𝑑𝑥 + (4𝑥 − 3𝑦 − 6) 𝑑𝑦 = 0

4. 𝑦(𝑥 3 𝑒 𝑥𝑦 − 𝑦) 𝑑𝑥 + 𝑥 (𝑦 + 𝑥 3𝑒 𝑥𝑦) 𝑑𝑦 = 0

5. 𝑦 2 𝑑𝑥 + (𝑥 2 − 𝑥𝑦 − 𝑦 2 ) 𝑑𝑦 = 0

6. (4𝑒 −𝑦 𝑠𝑖𝑛 𝑥 − 1) 𝑑𝑥 − 𝑑𝑦 = 0

7. 𝑥 4 𝑑𝑦 𝑑𝑥 + 𝑥 3𝑦 − 𝑠𝑒𝑐(𝑥𝑦) = 0

8. (1 + 𝑠𝑖𝑛 𝑦) 𝑑𝑥 𝑑𝑦 = 2𝑦 𝑐𝑜𝑠 𝑦 − 𝑥(𝑠𝑒𝑐 𝑦 + 𝑡𝑎𝑛 𝑦) 


1
Expert's answer
2021-08-30T16:52:48-0400

QUESTION 1



"\tan x\cdot\sin^2ydx+\cos^2x\cdot\cot ydy=0\longrightarrow\\[0.3cm]
\left.-\frac{\sin x}{\cos x}\cdot\sin^2ydx=\cos^2x\cdot\frac{\cos y}{\sin y}dy\right|\div\left(\sin^2y\cos^2x\right)\\[0.3cm]
-\frac{\sin xdx}{\cos^3x}=\frac{\cos ydy}{\sin^3y}\longrightarrow
\left[\begin{array}{l}
-\sin xdx=d\left(\cos x\right)\\[0.3cm]
\cos ydy=d\left(\sin y\right)
\end{array}\right]\\[0.3cm]
\int\frac{d\left(\cos x\right)}{\cos^3 x}=\int\frac{d\left(\sin y\right)}{\sin^3 y}\longrightarrow\\[0.3cm]
\left.-\frac{1}{2\cdot\cos^2x}=-\frac{1}{2\cdot\sin^2y}-\frac{Const}{2}\right|\cdot(-2)\\[0.3cm]
\boxed{\frac{1}{\cos^2x}=\frac{1}{\sin^2y}+Const}"

ANSWER



"\tan x\cdot\sin^2ydx+\cos^2x\cdot\cot ydy=0\longrightarrow\\[0.3cm]
\frac{1}{\cos^2x}=\frac{1}{\sin^2y}+Const-\text{implicit function}"

QUESTION 6



"\left(4e^{-y}\sin x-1\right)dx-dy=0\longrightarrow\\[0.3cm]
\frac{dy}{dx}=4e^{-y}\sin x-1\longrightarrow\boxed{y(x)=y_{hom.}(x)+y_{nonhom.}(x)}\\[0.3cm]
y_{hom.}(x) :\quad\left.\frac{dy}{dx}=4e^{-y}\sin x \right|\cdot\left(dx\cdot e^{y}\right)\\[0.3cm]
\int e^{y}dy=\int4\sin xdx\longrightarrow e^{y}=-4\cos x +C_1\longrightarrow\\[0.3cm]
\boxed{y_{hom.}(x)=\ln|C_1-4\cos x|}\\[0.3cm]
y_{nonhom.}(x) :\quad \frac{dy}{dx}=-1\longrightarrow\boxed{y_{nonhom.}(x)=-x}\\[0.3cm]
y(x)=y_{hom.}(x)+y_{nonhom.}(x)=\ln|C_1-4\cos x|-x\\[0.3cm]
\boxed{y(x)=\ln|C_1-4\cos x|-x}"

ANSWER



"\left(4e^{-y}\sin x-1\right)dx-dy=0\longrightarrow\\[0.3cm]
y(x)=\ln|C_1-4\cos x|-x"

QUESTION 3



"\left.\left(x-2y+1\right)dx+\left(4x-3y-6\right)dy=0\right|\div\left(\left(4x-3y-6\right)dx\right)\\[0.3cm]
\frac{x-2y+1}{4x-3y-6}+\frac{dy}{dx}=0\\[0.3cm]
\text{Let}\quad\left\{\begin{array}{l}
x=X+a\longrightarrow dx=dX\\[0.3cm]
y=Y+b\longrightarrow dy=dY
\end{array}\right.\quad\text{, where a,b - const}\\[0.3cm]
\frac{X+a-2\left(Y+b\right)+1}{4\left(X+a\right)-3\left(Y+b\right)-6}+\frac{dY}{dX}=0\\[0.3cm]
\frac{X-2Y+\left(a-2b+1\right)}{4X-3Y+\left(4a-3b-6\right)}+\frac{dY}{dX}=0\\[0.3cm]
\text{Let's choose constants "a" and "b" so}\quad\left\{\begin{array}{l}
a-2b+1=0\\[0.3cm]
4a-3b-6=0\end{array}\right.\\[0.3cm]
a-2b+1=4a-3b-6\longrightarrow3b-2b=4a-a-6-1\\[0.3cm]
\boxed{b=3a-7}\longrightarrow\boxed{a=3\quad b=2}\\[0.3cm]
\text{The choice was made to "a" and "b" were intact}\\[0.3cm]
\text{It remains to solve such an equation} : \frac{X-2Y}{4X-3Y}+\frac{dY}{dX}=0\\[0.3cm]
\text{Let}\quad Y=VX\longrightarrow\frac{dY}{dX}=V+X\frac{dV}{dX}\longrightarrow\\[0.3cm]
\frac{X-2VX}{4X-3VX}+V+X\frac{dV}{dX}=0\longrightarrow\\[0.3cm]
-X\frac{dV}{dX}=V+\frac{\cancel{X}\left(1-2V\right)}{\cancel{X}\left(4-3V\right)}\longrightarrow\\[0.3cm]
-X\frac{dV}{dX}=\frac{V\left(4-3V\right)+1-2V}{4-3V}\longrightarrow\\[0.3cm]
\left.-X\frac{dV}{dX}=\frac{-3V^2+2V+1}{4-3V}\right|\cdot\left(-\frac{\left(4-3V\right)dX}{X\left(-3V^2+2V+1\right)}\right)\longrightarrow\\[0.3cm]
\frac{\left(4-3V\right)dV}{3V^2-2V-1}=\frac{dX}{X}\\[0.3cm]
3V^2-2V-1=0\longrightarrow\sqrt{D}=\sqrt{\left(-2\right)^2-4\cdot3\cdot\left(-1\right)}=4\\[0.3cm]
V_1=\frac{2+4}{6}=1\quad\text{and}\quad V_2=\frac{2-4}{6}=-\frac{1}{3}\\[0.3cm]
3V^2-2V-1=3\left(V-1\right)\left(V+\frac{1}{3}\right)\equiv\left(V-1\right)\left(3V+1\right)\\[0.3cm]
4-3V=p\left(V-1\right)+q\left(3V+1\right)\longrightarrow\\[0.3cm]
4-3V=(p+3q)V+(q-p)\longrightarrow\\[0.3cm]
\left\{\begin{array}{l}
q-p=4\longrightarrow q=p+4\\[0.3cm]
p+3(p+4)=-3\longrightarrow 4p=-15
\end{array}\right.\longrightarrow\\[0.3cm]
\boxed{p=-\frac{15}{4}\quad\text{and}\quad q=\frac{1}{4}}\\[0.3cm]
\int\frac{\left(-\displaystyle\frac{15}{4}\left(V-1\right)+\displaystyle\frac{1}{4}\left(3V+1\right)\right)dV}{\left(V-1\right)\left(3V+1\right)}=\int\frac{dX}{X}\\[0.3cm]
\int\left(-\frac{15}{4}\cdot\frac{1}{3V+1}+\frac{1}{4}\cdot\frac{1}{V-1}\right)dV=\int\frac{dX}{X}\\[0.3cm]
\left.\frac{1}{4}\cdot\ln|V-1|-\frac{5}{4}\cdot\ln|3V+1|-\ln|X|=\ln|C|\right|\cdot\left(-4\right)\\[0.3cm]
5\ln|3V+1|+4\ln|X|-\ln|V-1|=-4\ln|C|\\[0.3cm]
\ln\left|\frac{\left(3V+1\right)^5X^4}{V-1}\right|=\ln\left|\frac{1}{C^4}\right|\equiv\ln|c|\\[0.3cm]
Y=VX\longrightarrow V=\frac{Y}{X}\longrightarrow\\[0.3cm]
\frac{\left(3\cdot\displaystyle\frac{Y}{X}+1\right)^5X^4}{\displaystyle\frac{Y}{X}-1}=c\longrightarrow\frac{\displaystyle\frac{\left(3Y+X\right)^5}{\cancel{X^5}}\cdot\cancel{X^4}}{\displaystyle\frac{Y-X}{\cancel{X}}}=c\\[0.3cm]
\frac{\left(3Y+X\right)^5}{Y-X}=c\longrightarrow\left\{\begin{array}{l}
x=X+3\longrightarrow X=x-3\\[0.3cm]
y=Y+2\longrightarrow Y=y-2
\end{array}\right.\\[0.3cm]
\frac{\left(3\left(y-2\right)+x-3\right)^5}{y-2-\left(x-3\right)}=c\longrightarrow
\boxed{\frac{\left(3y+x-9\right)^5}{y-x+1}=c}"

ANSWER



"\left(x-2y+1\right)dx+\left(4x-3y+6\right)dy=0\longrightarrow\\[0.3cm]
\frac{\left(3y+x-9\right)^5}{y-x+1}=c"

QUESTION 2



"\left.\left(1-2e^{x/y}\right)dx-2e^{x/y}\cdot\left(1-\frac{x}{y}\right)dy=0\right|\div\left(dy\right)\\[0.3cm]
\left(1-2e^{x/y}\right)\cdot\frac{dx}{dy}-2e^{x/y}\cdot\left(1-\frac{x}{y}\right)=0\\[0.3cm]
\text{Let}\quad x=yv\longrightarrow\frac{dx}{dy}=v+y\cdot\frac{dv}{dy}\\[0.3cm]
\left.\left(1-2e^v\right)\cdot\left(v+y\cdot\frac{dv}{dy}\right)-2e^v\cdot\left(1-v\right)=0\right|\div\left(1-2e^v\right)\\[0.3cm]
v+y\cdot\frac{dv}{dy}=\frac{2e^v-2ve^v}{1-2e^v}\longrightarrow y\cdot\frac{dv}{dy}=\frac{2e^v-2ve^v}{1-2e^v}-v\\[0.3cm]
y\cdot\frac{dv}{dy}=\frac{2e^v-2ve^v-v+2ve^v}{1-2e^v}\equiv\frac{2e^v-v}{1-2e^v}\\[0.3cm]
\left.y\cdot\frac{dv}{dy}=\frac{2e^v-v}{1-2e^v}\right|\cdot\left(\frac{dy}{y}\cdot\frac{1-2e^v}{2e^v-v}\right)\\[0.3cm]\\[0.3cm]
\int\frac{1-2e^v}{2e^v-v}dv=\int\frac{dy}{y}\longrightarrow\int\frac{d\left(v-2e^v\right)}{2e^v-v}=\int\frac{dy}{y}\\[0.3cm]
-\ln\left|v-2e^v\right|=\ln|y|+\ln|C|\longrightarrow\\[0.3cm]
\left.-\ln\left|y\cdot\left(v-2e^v\right)\right|=\ln|C|\right|\cdot\left(-1\right)\\[0.3cm]
\ln\left|y\cdot\left(\frac{x}{y}-2e^{x/y}\right)\right|=\ln\left|\frac{1}{C}\right|\equiv\ln|c|\\[0.3cm]
\boxed{x-2ye^{x/y}=c}"

ANSWER



"\left(1-2e^{x/y}\right)dx-2e^{x/y}\cdot\left(1-\frac{x}{y}\right)dy=0\\[0.3cm]
x-2ye^{x/y}=c"

QUESTION 4



"𝑦\left(𝑥^3\cdot 𝑒^{𝑥𝑦} − 𝑦\right)𝑑𝑥+𝑥\left(𝑦+𝑥^3\cdot 𝑒^{𝑥𝑦}\right)𝑑𝑦=0"

Let us check whether this equation in total differentials. Suppose,


"\underbrace{𝑦\left(𝑥^3\cdot 𝑒^{𝑥𝑦} − 𝑦\right)}_{M(x,y)}𝑑𝑥+\underbrace{𝑥\left(𝑦+𝑥^3\cdot 𝑒^{𝑥𝑦}\right)}_{N(x,y)}𝑑𝑦=0\\[0.3cm]
\frac{\partial M}{\partial y}=\frac{\partial N}{\partial x}-\text{condition must be fulfilled}\\[0.3cm]
\frac{\partial M}{\partial y}=\frac{\partial}{\partial y}\left(𝑦\left(𝑥^3\cdot 𝑒^{𝑥𝑦} − 𝑦\right)\right)=
x^3\cdot e^{xy}-y+𝑦\left(𝑥^4\cdot 𝑒^{𝑥𝑦} − 1\right)\\[0.3cm]
\boxed{\frac{\partial M}{\partial y}=x^3\cdot e^{xy}-y+𝑦𝑥^4\cdot 𝑒^{𝑥𝑦} − y}\\[0.3cm]
\frac{\partial N}{\partial x}=\frac{\partial}{\partial x}\left(x\left(y+𝑥^3\cdot 𝑒^{𝑥𝑦}\right)\right)=
y+𝑥^3\cdot 𝑒^{𝑥𝑦}+x\left(3x^2e^{xy}+x^3ye^{xy}\right)\\[0.3cm]
\boxed{\frac{\partial N}{\partial x}=y+x^3\cdot e^{xy}+3x^3e^{xy}+yx^4\cdot e^{xy}}\\[0.3cm]"

As we can see,



"x^3\cdot e^{xy}+yx^4\cdot e^{xy}-2y\neq y+4x^3\cdot e^{xy}+yx^4\cdot e^{xy}"

But,



"\frac{\partial M/\partial y-\partial N/\partial x}{N}=\\[0.3cm]
=\frac{x^3\cdot e^{xy}+yx^4\cdot e^{xy}-2y-y-4x^3\cdot e^{xy}-yx^4\cdot e^{xy}}{x\left(y+x^3\cdot e^{xy}\right)}=\\[0.3cm]
=\frac{-3\left(y+x^3\cdot e^{xy}\right)}{x\left(y+x^3\cdot e^{xy}\right)}=-\frac{3}{x}-\text{depends only on ''x''}"

This means that we can find the integrating factor :



"\int\frac{d\mu}{\mu}=\int\frac{-3dx}{x}\longrightarrow\ln|\mu|=-3\ln|x|\longrightarrow
\boxed{\mu=\frac{1}{x^3}}"

Then,



"\left.y\left(x^3\cdot e^{xy}-y\right)dx+x\left(y+x^3\cdot e^{xy}\right)dy=0\right|\cdot\left(\frac{1}{x^3}\right)\\[0.3cm]
\underbrace{\frac{y\left(x^3\cdot e^{xy}-y\right)}{x^3}}_{F_x(x,y)}\cdot dx+\underbrace{\frac{x\left(y+x^3\cdot e^{xy}\right)}{x^3}}_{F_y(x,y)}\cdot dy=0\\[0.3cm]
F_y(x,y)\equiv\frac{\partial F}{\partial y}=\frac{y+x^3\cdot e^{xy}}{x^2}\longrightarrow\\[0.3cm]
F(x,y)=\int\frac{y+x^3\cdot e^{xy}}{x^2}\cdot dy+f(x)\longrightarrow\\[0.3cm]
F(x,y)=\frac{1}{x^2}\cdot\left(\frac{y^2}{2}+\frac{x^3}{x}\cdot e^{xy}\right)+f(x)\longrightarrow\\[0.3cm]
\boxed{F(x,y)=\frac{y^2}{2x^2}+e^{xy}+f(x)}\\[0.3cm]
F_x(x,y)\equiv\frac{\partial F}{\partial x}=\frac{\partial}{\partial x}\left(\frac{y^2}{2x^2}+e^{xy}+f(x)\right)=\\[0.3cm]
-\frac{y^2}{x^3}+y\cdot e^{xy}+\frac{df}{dx}=\underbrace{\frac{y\left(x^3\cdot e^{xy}-y\right)}{x^3}}_{\text{from the equation}}\\[0.3cm]
-\frac{y^2}{x^3}+y\cdot e^{xy}+\frac{df}{dx}=y\cdot e^{xy}-\frac{y^2}{x^3}\longrightarrow\boxed{\frac{df}{dx}=0}\\[0.3cm]
\boxed{f(x)=Const}"

Conclusion,



"𝑦\left(𝑥^3\cdot 𝑒^{𝑥𝑦} − 𝑦\right)𝑑𝑥+𝑥\left(𝑦+𝑥^3\cdot 𝑒^{𝑥𝑦}\right)𝑑𝑦=0\longrightarrow\\[0.3cm]
d\left(\frac{y^2}{2x^2}+e^{xy}+Const\right)=0\longrightarrow\\[0.3cm]
\frac{y^2}{2x^2}+e^{xy}+Const=Const_1\longrightarrow\boxed{\frac{y^2}{2x^2}+e^{xy}=c}"

ANSWER



"𝑦\left(𝑥^3\cdot 𝑒^{𝑥𝑦} − 𝑦\right)𝑑𝑥+𝑥\left(𝑦+𝑥^3\cdot 𝑒^{𝑥𝑦}\right)𝑑𝑦=0\longrightarrow\\[0.3cm]
\frac{y^2}{2x^2}+e^{xy}=c"

QUESTION 5



"y^2dx+\left(x^2-xy-y^2\right)dy=0\longrightarrow\\[0.3cm]
y^2dx=-x^2\cdot\left(1-\frac{y}{x}-\frac{y^2}{x^2}\right)dy\longrightarrow\\[0.3cm]
\frac{dy}{dx}=\frac{-\displaystyle\frac{y^2}{x^2}}{1-\displaystyle\frac{y}{x}-\displaystyle\frac{y^2}{x^2}}\longrightarrow
\left\{\begin{array}{l}
y=xv\\[0.3cm]
\displaystyle\frac{dy}{dx}=v+x\cdot\displaystyle\frac{dv}{dx}
\end{array}\right.\\[0.3cm]
v+x\cdot\frac{dv}{dx}=\frac{-v^2}{1-v-v^2}\longrightarrow\\[0.3cm]
x\cdot\frac{dv}{dx}=\frac{-v\left(1-v-v^2\right)-v^2}{1-v-v^2}=\frac{-v+v^2+v^3-v^2}{1-v-v^2}\\[0.3cm]
\left.x\cdot\frac{dv}{dx}=\frac{v^3-v}{1-v-v^2}\right|\cdot\left(\frac{\left(1-v-v^2\right)dx}{x\left(v^3-v\right)}\right)\\[0.3cm]
\int\frac{\left(1-v^2\right)-v}{v\left(v^2-1\right)}dv=\int\frac{dx}{x}\longrightarrow\\[0.3cm]
\int\left(-\frac{1}{v}-\frac{1}{2}\cdot\frac{(v+1)-(v-1)}{(v-1)(v+1)}\right)dv=\ln|x|+\ln|C|\\[0.3cm]
\int\left(-\frac{1}{v}-\frac{1}{2}\cdot\frac{1}{(v-1)}+\frac{1}{2}\cdot\frac{1}{(v+1)}\right)dv=\ln|x|+\ln|C|\\[0.3cm]
-\ln|v|-\frac{1}{2}\ln|v-1|+\frac{1}{2}\cdot\ln|v+1|=\ln|Cx|\\[0.3cm]
\ln\left|\frac{1}{v}\cdot\sqrt{\frac{v+1}{v-1}}\right|=\ln|Cx|=\left[v=\frac{y}{x}\right]\\[0.3cm]
\frac{x}{y}\cdot\sqrt{\frac{\displaystyle\frac{y}{x}+1}{\displaystyle\frac{y}{x}-1}}=Cx\longrightarrow\boxed{\frac{x}{y}\cdot\sqrt{\frac{y+x}{y-x}}=Cx}"

ANSWER



"y^2dx+\left(x^2-xy-y^2\right)dy=0\longrightarrow\\[0.3cm]
\frac{x}{y}\cdot\sqrt{\frac{y+x}{y-x}}=Cx"

QUESTION 7



"x^4\cdot y'+x^3y=-\sec(xy)\longrightarrow\\[0.3cm]
\text{Let's make a replacement}\,\,\,u=xy\longrightarrow y=\frac{u}{x}\longrightarrow\boxed{ y'=\frac{x\cdot u'-u}{x^2}}\\[0.3cm]"

We substitute the replacement and the found derivative into the initial equation



"x^4\cdot\frac{x\cdot u'-u}{x^2}+x^3\cdot\frac{u}{x}=-\frac{1}{\cos u}\longrightarrow\\[0.3cm]
x^2\cdot\left(x\cdot u'-u+u\right)=-\frac{1}{\cos u}\longrightarrow\\[0.3cm]
\left.x^3\cdot\frac{du}{dx}=-\frac{1}{\cos u}\right|\cdot\left(\frac{\cos udx}{x^3}\right)\\[0.3cm]
\int\cos udu=\int\left(-\frac{dx}{x^3}\right)\longrightarrow\sin u=\frac{1}{2x^2}+c\longrightarrow\\[0.3cm]
u\equiv xy=\arcsin\left(c+\frac{1}{2x^2}\right)\longrightarrow\\[0.3cm]
\boxed{y=\frac{1}{x}\cdot\arcsin\left(c+\frac{1}{2x^2}\right)}"

ANSWER



"x^4\cdot y'+x^3y=-\sec(xy)\longrightarrow\\[0.3cm]
y=\frac{1}{x}\cdot\arcsin\left(c+\frac{1}{2x^2}\right)"

QUESTION 8



"\left(1+\sin 𝑦\right)\cdot\frac{dx}{dy}=2𝑦\cos 𝑦−𝑥\left(\sec 𝑦+\tan 𝑦\right)\longrightarrow\\[0.3cm]

y_{hom.}(x) :\quad\left(1+\sin 𝑦\right)\cdot\frac{dx}{dy}=-x\left(\frac{1}{\cos y}+\frac{\sin y}{\cos y}\right)\\[0.3cm]
\left.\left(1+\sin 𝑦\right)\cdot\frac{dx}{dy}=-x\left(\frac{1+\sin y}{\cos y}\right)\right|\cdot\left(\frac{dy}{x\left(1+\sin y\right)}\right)\\[0.3cm]
\int\frac{dx}{x}=-\int\frac{dy}{\cos y}\longrightarrow\ln|x|=-\int\frac{\cos y dy}{\cos^2y}\boxed{\longrightarrow}\\[0.3cm]
\int\frac{\cos ydy}{\cos^2 y}=\int\frac{d\left(\sin y\right)}{1-\sin^2 y}=\\[0.3cm]
=\int\frac{1}{2}\cdot\left(\frac{(1+\sin y)+(1-\sin y)}{(1-\sin y)(1+\sin y)}\right)d\left(\sin y\right)=\\[0.3cm]
=\frac{1}{2}\cdot\int\left(\frac{1}{1-\sin y}+\frac{1}{1+\sin y}\right)d\left(\sin y\right)=\\[0.3cm]
=\frac{1}{2}\cdot\left(-\ln|1-\sin y|+\ln|1+\sin y|\right)=\\[0.3cm]
=\frac{1}{2}\cdot\ln\left|\frac{1+\sin y}{1-\sin y}\right|=\ln\left|\sqrt{\frac{\left(1+\sin y\right)^2}{(1-\sin y)(1+\sin y)}}\right|=\\[0.3cm]
=\ln\left|\frac{1+\sin y}{\sqrt{1-\sin^2y}}\right|=\ln\left|\frac{1+\sin y}{\cos y}\right|\\[0.3cm]
\boxed{\longrightarrow}\ln|x|=\ln|C|-\ln\left|\frac{1+\sin y}{\cos y}\right|\longrightarrow\\[0.3cm]
\ln|x|=\ln\left|\frac{C\cos y}{1+\sin y}\right|\longrightarrow\boxed{x(y)=\frac{C\cos y}{1+\sin y}}"



Let's apply the constant variation method "C\longrightarrow C(y) :"



"\frac{dx}{dy}=\frac{\left(C'\cos y-C\sin y\right)\left(1+\sin y\right)-C\cos^2y}{\left(1+\sin y\right)^2}\\[0.3cm]
\left(1+\sin y\right)\cdot\frac{dx}{dy}=2y\cos y-x\left(\frac{1+\sin y}{\cos y}\right)\\[0.3cm]
\left(1+\sin y\right)\cdot\frac{\left(C'\cos y-C\sin y\right)\left(1+\sin y\right)-C\cos^2y}{\left(1+\sin y\right)^2}=\\[0.3cm]
=2y\cos y-\frac{C\cos y}{1+\sin y}\cdot\frac{1+\sin y}{\cos y}\\[0.3cm]
\frac{C'\cos y\left(1+\sin y\right)-C\sin y-C\sin^2y-C\cos^2}{1+\sin y}=2y\cos y-C\\[0.3cm]
C'\cos y-C\cdot\frac{\sin y+\overbrace{\left(\sin^2 y+\cos^2 y\right)}^{=1}}{1+\sin y}=2y\cos y-C\\[0.3cm]
C'\cos y-\cancel{C}\cdot\frac{\cancel{1+\sin y}}{\cancel{1+\sin y}}=2y\cos y-\cancel{C}\\[0.3cm]
\left.\frac{dC}{dy}\cdot\cancel{\cos y}=2y\cdot\cancel{\cos y}\right|\cdot dy\\[0.3cm]
\int dC=\int2ydy\longrightarrow\boxed{C(y)=y^2+c}"

Conclusion,



"x(y)=\frac{C(y)\cdot\cos y}{1+\sin y}\equiv\frac{\left(y^2+c\right)\cdot\cos y}{1+\sin y}"

ANSWER



"\left(1+\sin 𝑦\right)\cdot\frac{dx}{dy}=2𝑦\cos 𝑦−𝑥\left(\sec 𝑦+\tan 𝑦\right)\longrightarrow\\[0.3cm]
x(y)=\frac{\left(y^2+c\right)\cdot\cos y}{1+\sin y}"

QUESTION ( first line equation )

I just repeat the solution to equation # 2.



"\left.\left(1-2e^{x/y}\right)dx-2e^{x/y}\cdot\left(1-\frac{x}{y}\right)dy=0\right|\div\left(dy\right)\\[0.3cm]
\left(1-2e^{x/y}\right)\cdot\frac{dx}{dy}-2e^{x/y}\cdot\left(1-\frac{x}{y}\right)=0\\[0.3cm]
\text{Let}\quad x=yv\longrightarrow\frac{dx}{dy}=v+y\cdot\frac{dv}{dy}\\[0.3cm]
\left.\left(1-2e^v\right)\cdot\left(v+y\cdot\frac{dv}{dy}\right)-2e^v\cdot\left(1-v\right)=0\right|\div\left(1-2e^v\right)\\[0.3cm]
v+y\cdot\frac{dv}{dy}=\frac{2e^v-2ve^v}{1-2e^v}\longrightarrow y\cdot\frac{dv}{dy}=\frac{2e^v-2ve^v}{1-2e^v}-v\\[0.3cm]
y\cdot\frac{dv}{dy}=\frac{2e^v-2ve^v-v+2ve^v}{1-2e^v}\equiv\frac{2e^v-v}{1-2e^v}\\[0.3cm]
\left.y\cdot\frac{dv}{dy}=\frac{2e^v-v}{1-2e^v}\right|\cdot\left(\frac{dy}{y}\cdot\frac{1-2e^v}{2e^v-v}\right)\\[0.3cm]\\[0.3cm]
\int\frac{1-2e^v}{2e^v-v}dv=\int\frac{dy}{y}\longrightarrow\int\frac{d\left(v-2e^v\right)}{2e^v-v}=\int\frac{dy}{y}\\[0.3cm]
-\ln\left|v-2e^v\right|=\ln|y|+\ln|C|\longrightarrow\\[0.3cm]
\left.-\ln\left|y\cdot\left(v-2e^v\right)\right|=\ln|C|\right|\cdot\left(-1\right)\\[0.3cm]
\ln\left|y\cdot\left(\frac{x}{y}-2e^{x/y}\right)\right|=\ln\left|\frac{1}{C}\right|\equiv\ln|c|\\[0.3cm]
\boxed{x-2ye^{x/y}=c}"

ANSWER



"\left(1-2e^{x/y}\right)dx-2e^{x/y}\cdot\left(1-\frac{x}{y}\right)dy=0\\[0.3cm]
x-2ye^{x/y}=c"